OpenStudy (anonymous):

Rewrite the equation 2x + 3y = 6 in slope-intercept form.

OpenStudy (anonymous):

get y by itself.

OpenStudy (johnweldon1993):

how would you get "y" by itself on that left side?

OpenStudy (tkhunny):

Subtract 2x. Divide by 3.

OpenStudy (anonymous):

i still don't get it

OpenStudy (tkhunny):

Okay, one at a time. Subtract 2x. You won't GET it until after you DO it.

OpenStudy (anonymous):

y=mx+b 3y=6-2x y=-1x+2 (this is slope intercept form)